OpenStudy (annie96):

The set {0,1} is closed under which operation? none of the above addition multiplication subtraction

OpenStudy (anonymous):

A set has closure if you can do that operation in any direction and always get a number that belongs to that set.

jimthompson5910 (jim_thompson5910):

A set is closed under an operation if you can take any 2 numbers in the set, perform the operation on them, and have the result be a number in the set For example: {1, 2, 3} this set is NOT closed under addition because 2+3 = 5. The numbers 2 and 3 are in the set, but 5 is not.

OpenStudy (anonymous):

lol nice

OpenStudy (anonymous):

So for this one: 0+1 = 1 But also 0 * 1 = 0 So... it should be multiplication and addition, but thats not a choice. @jim_thompson5910 thoughts?

jimthompson5910 (jim_thompson5910):

addition won't work because 1+1 = 2 is not in the set

jimthompson5910 (jim_thompson5910):

same idea for subtraction

OpenStudy (anonymous):

oh ok, sorry I didn't think that way. so its going to be multiplication

jimthompson5910 (jim_thompson5910):

yes 0*0 = 0 0*1 = 0 1*0 = 0 1*1 = 1 a table is usually a good way to sort it all out